# Adaptive operator learning for infinite-dimensional Bayesian inverse problems

### v2

@@ 4.3 Example 3: The reaction diffusion problem
We only take into account the OOD data as Example 1 for the inverse problem. In other words, we will attempt to inverse the first 128 KL modes using the ground truth $\mathbf{m}_{ref}(\mathbf{x})$, which is defined by (33) with $\theta_{k}\sim\mathcal{U}[-20,20],k=1,\cdots,256$. The exact solution and the corresponding synthetic data are displayed in Fig.15.
@@
Fig. 15: The reaction diffusion case. Left: the true initial field $m_{ref}(\mathbf{x})$. Right: the pressure field $u(\mathbf{x})$ and the corresponding $36$ equidistant observations with noise level 0.01.
The numerical results are shown in Fig.16. We can clearly see that refinement reduces the local model error significantly, and thus the inversion error will continue to decrease. This implies that our surrogate model can maintain its local accuracy during the inversion process by focusing on the region with the highest posterior probability. Finally, after six iterations of the initial model, the retraining was terminated using the stop criteria. Furthermore, as shown in the right display of Fig.16, DeepOnet-UKI-Adaptive can achieve nearly the same order of accuracy as FEM-UKI. This can be further confirmed by examining at Figs.17 and 18, which plot the final estimated initial fields and estimated states obtained by different methods.
@@
Fig. 16: The inversion results obtained by three methods for the reaction diffusion case. Left: the data fitting error. Middle: the model error. Right: the relative inversion errors.
@@
Fig. 17: The inversion results for reaction diffusion case. Above: the estimated initial fields. Below: the absolute errors with respect to the true ones.
@@
Fig. 18: The inversion results for the reaction diffusion case. Above: the approximate states evaluated at the final estimated initial fields. Below: the absolute error between the approximate state and the true state.
@@
Fig. 19: Left: the error box plot of $e_{\mathcal{I}-DeepOnet}-e_{\mathcal{I}-FEM}$. Right: the mean total number of forward evaluations for FEM-UKI and DeepOnet-UKI-Adaptive.
@@ 5 Conclusion
We present an adaptive operator learning framework for iteratively reducing the model error in Bayesian inverse problems. In particular, the unscented Kalman inversion(UKI) is used to approximate the solution of inverse problems, and the DeepOnet is utilized to construct the surrogate. We suggest a greedy algorithm to choose the adaptive samples to retrain the approximate model. The performance of the proposed strategy has been illustrated by three numerical examples. Although only the UKI algorithm is considered in this paper, the framework can be conveniently extended to a much wider class of particle-based methods with simple and minor modifications. The extension of the present algorithm with other neural operators is also straightforward.
@@
Note that by Eq.(3.4), we have

### v3

@@ 4.3 Example 3: The reaction diffusion problem
We only take into account the OOD data as Example 1 for the inverse problem. In other words, we will attempt to inverse the first 128 KL modes using the ground truth $\mathbf{m}_{ref}(\mathbf{x})$, which is defined by (33) with ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\zeta_{k}}\sim\mathcal{U}[-20,20],k=1,\cdots,256$. The exact solution and the corresponding synthetic data are displayed in Fig.18.
@@
Figure 18: The reaction diffusion case. Left: the true initial field $m_{ref}(\mathbf{x})$. Right: the pressure field $u(\mathbf{x})$ and the corresponding $36$ equidistant observations with noise level 0.01.
The numerical results are shown in Fig.19. We can clearly see that refinement reduces the local model error significantly, and thus the inversion error will continue to decrease. This implies that our surrogate model can maintain its local accuracy during the inversion process by focusing on the region with the highest posterior probability. Finally, after six iterations of the initial model, the retraining was terminated using the stop criteria. Furthermore, as shown in the right display of Fig.19, DeepOnet-UKI-Adaptive can achieve nearly the same order of accuracy as FEM-UKI. This can be further confirmed by examining at Figs.20 and 21, which plot the final estimated initial fields and estimated states obtained by different methods. In this case, the CPU time of evaluating the conventional FEM-UKI is more than 5149s. In contrast, for the DeepOnet-UKI-Adaptive approach, the online CPU times is only 530s, meaning that the adaptive approach can provide accurate results, yet with less computational time.
@@
Figure 19: The inversion results obtained by three methods for the reaction diffusion case. Left: the data fitting error. Middle: the model error. Right: the relative inversion errors.
[…]
Figure 20: The inversion results for reaction diffusion case. Above: the estimated initial fields. Below: the absolute errors with respect to the true ones.
[…]
Figure 21: The inversion results for the reaction diffusion case. Above: the approximate states evaluated at the final estimated initial fields. Below: the absolute error between the approximate state and the true state.
[…]
Figure 22: Left: the error box plot of $e_{\mathcal{I}-DeepOnet}-e_{\mathcal{I}-FEM}$. Right: the mean total number of forward evaluations for FEM-UKI and DeepOnet-UKI-Adaptive.
@@ 5 Conclusion
We present an adaptive operator learning framework for iteratively reducing the model error in Bayesian inverse problems. In particular, the unscented Kalman inversion(UKI) is used to approximate the solution of inverse problems, and the DeepOnet is utilized to construct the surrogate. We suggest a greedy algorithm to choose the adaptive samples to retrain the approximate model. The performance of the proposed strategy has been illustrated by three numerical examples. It should be noted that our adaptive framework may be ineffective in certain situations due to the use of Kalman-based methods. On the one hand, for strictly non-Gaussian posterior distributions, such as those with multiple modes, the UKI fails to capture all of the modes because it approximates the posterior using a Gaussian distribution. On the other hand, for some low- to moderate-dimensional problems, particularly when focusing on a specific task, the CES framework may be more suitable. Nonetheless, our strategy is intended to be adaptable to a broader range of scenarios. Using this framework, we can easily modify the posterior computation and surrogate methods to broaden the applicability of our method. While future work will address the potential drawbacks mentioned earlier.
@@
Note that by Eq.(18), we have
@@
Acknowledgment
The authors would like to thank anonymous referees for their many insightful and constructive comments and suggestions that improved the organization and quality of our paper essentially.